OpenStudy (anonymous):

finnd the solution to the system of equations use commas to seperate the x- and y-values. x= 0 +y 6x-5y=23

OpenStudy (lynfran):

\[x=0+y\]\[6x-5y=23\] \[6x=23+5y\]\[x=\frac{ 23+5y }{ 6 }\] now sub \[x=\frac{ 23+5y }{ 6 }\]in \[x=0+y\]we get\[\frac{ 23+5y }{ 6 }=0+y\]multiple both side of the equation by 6..we get\[23+5y=6y\]now we solve for y..\[23=6y-5y\]\[y=23\]now since we find y...we can find x..by by substituting y=23 in the 1st or 2nd equation...so, using equation 1\[x=0+y\]\[x=0+23\]\[x=23\]

OpenStudy (anonymous):

thanks

OpenStudy (lynfran):

welcome please medal thanks